Use Green’s theorem to evaluate ∫ (x^5+2y)dx + (4x-6y^3)dy where C is the circle x^2 + y ^2 = 4 .
1
Expert's answer
2019-10-07T10:08:35-0400

ANSWER¨ 8π

EXPLANATION.Let "D=\left\{ \left( x,y \right) :\quad { x }^{ 2 }+{ y }^{ 2 }\le 4 \right\}" . The area of the D is 4π="\iint_{D} dxdy" By the Green's theorem "\oint { \left( { x }^{ 5 }+2y \right) dx+\left( 4x-6{ y }^{ 3 } \right) dy\quad } =" "\iint_{D} \left [ \frac{\partial (4x-6y^{3})}{\partial x } -\frac{\partial(x^{5}+2y) }{\partial y}\right ]dxdy" ="\iint_{D} \left [ 4-2\right ]dxdy" =8π.